SPARO: Recent Progress
On Oct. 12th, 1999, SPARO was installed on Viper by Austral Winter 1999 winter-overs
Griffin and John Davis. The installation had been planned
for July, but had to be postponed due a failure of Viper's azimuth motor that rendered
the telescope inoperable for several months. During the remaining few
weeks of the winter season,
successful cryogenic operation of SPARO was demonstrated: Helium-4
and Helium-3 hold times were five days and four days, respectively. Other system
tests were successfully carried out, including measurements of amplifier noise
due to the JFET amplifiers in the SPARO cryostat. Due to a computer malfunction
and other minor setbacks, no observations were possible before
Novak, Chuss, and Matt Newcomb of Carnegie-Mellon arrived at Pole on Nov. 3,
1999. After some days of debugging computer problems, "first-light at
Pole" was obtained on Nov.
11, by observing the Moon. SPARO was operated through the third week of November.
Tests that were carried out during this SPARO run include
(1) Pointing: Both the Moon and the bright Galactic cloud
RCW-57 were found to be within a few tenths of a degree of the nomimal Viper
(2) Beam-size: The beam's FWHM was found to be approximately 4-6
arcminutes, close to the nominal size of 4 arcminutes.
(3) Signal strength:
Observations of the Moon indicate that the coupling of SPARO's beam to the sky is
reasonably high - the Moon signal was approximately equal to the expected value.
Further analysis is underway and should produce an accurate measurement of the
optical efficiency of the instrument-plus-telescope system.
offsets: These were found to be extremely small, as expected for this
under-illuminated optical design.
(5) Noise: the mechanical pump that sits out
on the telescope and pumps on SPARO's 1.5 K stage was found to be causing
occasional microphonics noise. To eliminate this problem, the pump will be moved
off the telescope for winter 2000 observations. We will pump through the new,
more rugged Helium pressure hoses that are currently being installed in Viper's
cable wrap for use with the Corona instrument. Aside from these occasional
microphonics due to the pump, we were at all times limited by fundamental
bolometer and photon noise, as expected.
(6) System operation: Data acquistion
and analysis operations were successfully and repeatedly tested for mapping,
photometric, and polarimetric modes.
We are now preparing for SPARO's 2000 winter-over by purchasing spares for
critical parts, and by refining SPARO's documentation and observing plan.
Renbarger is analysing the test results from Nov. 1999. Chuss and Novak will be
returning to Pole in late January to cool SPARO and reinstall it onto the
Greg Griffin of Carnegie-Mellon University and Dave Pernic of the University of
Chicago's Yerkes Observatory will operate SPARO during Winter 2000. Observations
are scheduled for February through May.